# Discontinuous Galerkin discretization of conservative dynamical low-rank approximation schemes for the Vlasov–Poisson equation

### v1

@@ 3.3 Conservation of mass and momentum
$\Delta\rho=\sum_{k}\tilde{X}^{n+1}_{k}\,$ $\Big{(}\sum_{ijs}\big{[}(\tilde{X}_{k}^{n+1},E_{s}^{n}X_{i}^{n})_{x}\,S_{ij}^{n}\,\langle\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),1\rangle_{\mathcal{T}_{h}^{(v)}}$
$-\langle\hat{\mathrm{d}}_{x_{s}}X_{i}^{n},\tilde{X}_{k}^{n+1}\rangle_{\mathcal{T}_{h}^{(x)}}\,S_{ij}^{n}\,(1,v_{s}V_{j}^{n})_{v,\omega}\big{]}\Big{)}.$
This follows by using $\ell=1$ in (37) and taking $V^{n}_{1}=\tilde{V}^{n+1}_{1}=1/\|1\|_{v,\omega}$ into account. Applying Lemma 3.1 shows that
$\langle\hat{\mathrm{d}}_{v_{s}}(\omega V_{j}^{n}),1\rangle_{\mathcal{T}_{h}^{(v)}}=0,$
and hence, the first term in the formula for $\Delta\rho$ vanishes. For the second term we use the Riesz representation $\hat{\Phi}_{X_{i}^{n},s}$ of $\hat{\mathrm{d}}_{x_{s}}X_{i}^{n}$, giving
$\Delta\rho=-\sum_{ijs}\,S_{ij}^{n}\,(1,v_{s}V_{j}^{n})_{v,\omega}\sum_{k}\tilde{X}^{n+1}_{k}\,\big{(}\hat{\Phi}_{X_{i}^{n},s},\tilde{X}^{n+1}_{k}\big{)}_{x}=-\sum_{ijs}\,S_{ij}^{n}\,(1,v_{s}V_{j}^{n})_{v,\omega}\hat{\Phi}_{X_{i}^{n},s},$
using that $\hat{\Phi}_{X_{i}^{n},s}$ is contained in augmented subspace spanned by the orthonormal $\tilde{X}^{n+1}_{i}$. Plugging this into (43), testing it with $\varphi\in\mathcal{V}_{h,p}^{(x)}$ and using the linearity of $\hat{\mathrm{d}}_{x_{s}}$ yields
$\frac{1}{\tau}(\rho^{n+1}-\rho^{n},\varphi)_{x}$ $=-\sum_{ijs}\,S_{ij}^{n}\,(1,v_{s}V_{j}^{n})_{v,\omega}\,(\hat{\Phi}_{X_{i}^{n},s},\varphi)_{x}$
$=-\sum_{ijs}\,S_{ij}^{n}\,(1,v_{s}V_{j}^{n})_{v,\omega}\,\langle\hat{\mathrm{d}}_{x_{s}}X_{i}^{n},\varphi\rangle_{\mathcal{T}_{h}^{(x)}}$
$=-\sum_{s}\langle\hat{\mathrm{d}}_{x_{s}}\big{(}\sum_{ij}X_{i}^{n}S_{ij}^{n}(V_{j}^{n},v_{s})_{v,\omega}\big{)},\varphi\rangle_{\mathcal{T}_{h}^{(x)}},$
which, when compared with (41), exactly means
$\frac{1}{\tau}(\rho^{n+1}-\rho^{n},\varphi)_{x}=-\sum_{s}\langle\hat{\mathrm{d}}_{s}j_{s}^{n},\varphi\rangle_{x}.$
Applying Lemma 3.1 yields (42). Letting $\varphi=1\in\mathcal{V}_{h,p}^{(x)}$ in (42) shows conservation of total mass. ∎
Next we consider conservation of momentum. In contrast to mass conservation, our proof for momentum conservation requires that the central flux, $\alpha=1$ in (31), is used, at least in the K-step. We will need the following lemma, which implies that the projection of $\bm{E}^{n}\rho^{n}$ onto the augmented basis $\tilde{\bm{X}}^{n+1}$ already yields its projection onto the entire DG space $\mathcal{V}_{h,p}^{(x)}$. Here $\bm{E}^{n}$ is the electric field computed from the data in the $n$-th step.
Lemma 3.3.
Assume a central flux, i.e. $\alpha=1$ in (31), for computing the K-steps (35). If $m\geq 1+d$, then for $n=0,1,2,\dots$, it holds that the $L_{2}$-orthogonal projection of $E^{n}_{s}\rho^{n}$ onto $\mathcal{V}_{h,p}^{(x)}$ is contained in the subspace spanned by the augmented basis $\tilde{\bm{X}}^{n+1}$.
Proof.
Applying Lemma 3.1 (integration by parts) to the K-step formula (35) gives
$\frac{1}{\tau}\big{[}(\bm{K}^{n+1},\bm{\psi})_{x}-(\bm{K}^{n},\bm{\psi})_{x}\big{]}$ $=-\sum_{s^{\prime}}\big{[}\langle\hat{\mathrm{d}}_{x_{s^{\prime}}}(\mathsf{A}^{(x,{s^{\prime}})}\cdot\bm{K}^{n}),\bm{\psi}\rangle_{x}-E^{n}_{s^{\prime}}(\mathsf{B}_{h}^{(x,s^{\prime})}\cdot\bm{K}^{n},\bm{\psi})_{x}\big{]}$
$+\sum_{e\in\mathcal{F}_{h}^{(x)}}\sum_{s^{\prime}}\int_{e}n^{(e)}_{s^{\prime}}\,(\mathsf{A}^{(x,s^{\prime})}\cdot\{\bm{K}^{n}\}_{e})\cdot[\bm{\psi}]_{e}\,\mathrm{d}S$
$-\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}\bm{F}^{*}_{e}((\bm{K}_{e}^{n})^{-},(\bm{K}_{e}^{n})^{+})\cdot[\bm{\psi}]_{e}\,\mathrm{d}S,$
where we also used the linearity of $\hat{\mathrm{d}}_{x_{s^{\prime}}}$. By definition of the numerical flux (31), the edge contributions cancel out for $\alpha=1$. Testing with function of the form $\bm{\psi}=\varphi\,\bm{e}_{1+s}$ (that is, $\bm{\Psi}$ has $\varphi\in\mathcal{V}_{h,p}^{(x)}$ at component $1+s$, the others are zero), where $s$ will be the component of $\bm{E}^{n}$ later, gives
$\frac{1}{\tau}\big{[}(K_{1+s}^{n+1},\varphi)_{x}-(K_{1+s}^{n},\varphi)_{x}\big{]}=-\sum_{s^{\prime}}\sum_{j}\big{[}\mathsf{A}^{(x,s^{\prime})}_{1+s,j}\cdot\langle\hat{\mathrm{d}}_{x_{s^{\prime}}}K^{n}_{j},\varphi\rangle_{x}-E^{n}_{s^{\prime}}\big{(}(\mathsf{B}^{(x,s^{\prime})}_{h})_{1+s,j}\cdot K_{j}^{n},\varphi)_{x}\big{]}.$ (44)
By the definition of the matrix elements (34) of $\mathsf{B}_{h}^{(x,s^{\prime})}$ and Lemma 3.1 it holds that
$(\mathsf{B}^{(x,s^{\prime})}_{h})_{1+s,j}=\langle\hat{\mathrm{d}}_{v_{s^{\prime}}}(\omega V_{j}^{n}),v_{s}/\|v_{s}\|_{v,\omega}\rangle_{v}=-\frac{1}{\|v_{s}\|_{v,\omega}}(V_{j}^{n},1)_{v,\omega}\delta_{s^{\prime}s}=-\frac{\|1\|_{v,\omega}}{\|v_{s}\|_{v,\omega}}\delta_{j1}\delta_{s^{\prime}s}.$
Inserting this in (44), using the expression for $\rho$ from (40), and rearranging one obtains
$\big{(}E_{s}^{n}\rho^{n},\varphi\big{)}_{x}=\big{(}P,\varphi\big{)}_{x}\quad\text{for all }\varphi\in\mathcal{V}_{h,p}^{(x)}$ (45)
with
$P\coloneqq\frac{\|v_{s}\|_{v,\omega}}{\tau}K_{1+s}^{n+1}-\frac{\|v_{s}\|_{v,\omega}}{\tau}K_{1+s}^{n}+\|v_{s}\|_{v,\omega}\sum_{s^{\prime},j}\mathsf{A}^{(x,s^{\prime})}_{1+s,j}\hat{\Phi}_{K^{n}_{j},s^{\prime}}.$
Note that $P$ is a linear combination of functions which by construction are contained in the span of the augmented basis $\tilde{\bm{X}}^{n+1}$, which is a subspace of $\mathcal{V}_{h,p}^{(x)}$. From (45), it follows that $P$ is the orthogonal projection of $E_{s}^{n}\rho^{n}$ onto $\mathcal{V}_{h,p}^{(x)}$. ∎
@@
$\frac{1}{\tau}(j_{s}^{n+1}-j_{s}^{n},\varphi)_{x}+\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}-\bm{\sigma}_{s}^{n}\cdot\nabla\varphi+E^{n}_{s}\rho^{n}\varphi\,\mathrm{d}\bm{x}+\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}\big{(}\bm{n}^{(e)}\cdot\{\bm{\sigma}_{s}^{n}\}_{e}\big{)}\,[\varphi]_{e}\,\mathrm{d}S=0$ (46)
for all $\varphi\in\mathcal{V}_{h,p}^{(x)}$. This corresponds to a DG discretization of the continuity equations (39) using central fluxes and explicit Euler time stepping as in (32). Hence, the discrete local continuity equation holds for the momentum density and the total momentum $\int_{\Omega^{(x)}}\bm{j}^{n}\,\mathrm{d}\bm{x}$ is conserved.
@@
Combining (41) with (36) and (37), the updated momentum densities can be expressed as
$j_{s}^{n+1}=\|v_{s}\|_{v,\omega}\sum_{k}\tilde{X}^{n+1}_{k}S^{n+1}_{k,1+s}=j_{s}^{n}+\tau\Delta j_{s}\$ (47)
where
$\Delta j_{s}=\sum_{k}\tilde{X}^{n+1}_{k}\,$ $\Big{(}\sum_{ijs^{\prime}}\big{[}(\tilde{X}_{k}^{n+1},E_{s^{\prime}}^{n}X_{i}^{n})_{x}\,S_{ij}^{n}\,\langle\hat{\mathrm{d}}_{v_{s^{\prime}}}(\omega V_{j}^{n}),v_{s}\rangle_{\mathcal{T}_{h}^{(v)}}$
$-\langle\hat{\mathrm{d}}_{x_{s^{\prime}}}X_{i}^{n},\tilde{X}_{k}^{n+1}\rangle_{\mathcal{T}_{h}^{(x)}}\,S_{ij}^{n}\,(v_{s},v_{s^{\prime}}V_{j}^{n})_{v,\omega}\big{]}\Big{)}$
This follows by using $\ell=1+s$ in (37) and taking $V^{n}_{1+s}=\tilde{V}^{n+1}_{1+s}=v_{s}/\|v_{s}\|_{v,\omega}$ into account. By Lemma 3.1, it holds
$\langle\hat{\mathrm{d}}_{v_{s^{\prime}}}(\omega V_{j}^{n}),v_{s}\rangle_{\mathcal{T}_{h}^{(v)}}=-(V_{j}^{n},1)_{v,\omega}\delta_{s^{\prime}s}=-\|1\|_{v,\omega}\delta_{j1}\delta_{s^{\prime}s}$
in the first term. For the second term we note that the Riesz representative $\hat{\Phi}_{X_{i}^{n},s^{\prime}}$ of $\hat{\mathrm{d}}_{x_{s^{\prime}}}X_{i}^{n}$ is contained in the span of the orthonormal basis $\tilde{X}_{k}^{n+1}$. As a result,
$\Delta j_{s}$ $=-\sum_{k}(\tilde{X}_{k}^{n+1},E_{s}^{n}\|1\|_{v,\omega}\sum_{i}X_{i}^{n}S_{i1}^{n})_{x}\,\tilde{X}_{k}^{n+1}-\sum_{s^{\prime}}(v_{s}v_{s^{\prime}},\sum_{ij}\omega S_{ij}^{n}V_{j}^{n})_{v}\hat{\Phi}_{X_{i}^{n},s^{\prime}}$
$=-\sum_{k}(\tilde{X}_{k}^{n+1},E_{s}^{n}\rho^{n})_{x}\,\tilde{X}_{k}^{n+1}-\sum_{s^{\prime}}(v_{s}v_{s^{\prime}},\sum_{ij}\omega S_{ij}^{n}V_{j}^{n})_{v}\hat{\Phi}_{X_{i}^{n},s^{\prime}}$
where for the second equality we used the formula (40) for $\rho$. Note that the first sum over $k$ equals the orthogonal projection of $E^{n}_{s}\rho^{n}$ onto the span of the orthogonal basis $\bm{\tilde{X}}^{n+1}$. By Lemma 3.3, this projection is equivalent to the full space projection onto $\mathcal{V}_{h,p}^{(x)}$. Therefore, testing this sum with $\varphi\in\mathcal{V}_{h,p}^{(x)}$ just gives $(E_{s}^{n}\rho^{n},\varphi)_{x}$. Also using the definition of $\bm{\sigma}_{s}$ from (39) we conclude
$(\Delta j_{s},\varphi)_{x}=-\big{(}E_{s}^{n}\rho^{n},\varphi\big{)}_{x}-\sum_{s^{\prime}}\langle\hat{\mathrm{d}}_{x_{s^{\prime}}}(\bm{\sigma}_{s}^{n})_{s^{\prime}},\varphi\rangle_{x}.$
Combining with (47) and applying Lemma 3.1, we obtain the desired equation (46).
For the conservation of momentum we note that in the continuous setting one has [10]
@@
This also holds for the discrete versions since the computed $\bm{E}^{n}$ is the exact electric field with respect to the (discrete) density $\rho^{n}$, see section 3.2. Setting $\varphi=1\in\mathcal{V}_{h,p}^{(x)}$ in (46), and using the fact that $\int_{\Omega^{(x)}}\bm{E}\,\mathrm{d}\bm{x}=0$, we conclude that the total momentum is conserved. ∎
@@ 3.4 Spatial adaptivity
We consider the general setup of section 3 (which applies to both the K- and L-step) and employ an elementwise error indicator based on the projection of a DG solution. Specifically, given a DG function $\bm{U}\in(V_{h,p}^{(y)})^{r}$ of polynomial order $p$, we project it onto a DG space of order $p-1$ within each element $T\in\mathcal{T}_{h}^{(y)}$, resulting in a function $\tilde{\bm{U}}\in V_{h,p-1}^{(y)}$ that satisfies
$(\tilde{\bm{U}},\varphi)_{L_{2}(\Omega^{(y)})}=(\bm{U},\varphi)_{L_{2}(\Omega^{(y)})}\quad\text{for all }\varphi\in\mathcal{V}_{h,p-1}^{(y)}.$
@@
In order to remove unnecessary refinements and maintain computational efficiency, a coarsening procedure is applied. The same error indicator is used to assess whether refinement is still required. Specifically, all children of a parent element are removed if the sum of their error indicators falls below $c\,\epsilon$, where $10<c<$ 1 is a safety factor.
@@ 4.1 Landau damping in one spatial dimension
The results for the conservative unconventional integrator with truncation to a fixed rank of $10$, shown in Figure 1, closely resemble those in [10, Fig. 1]. While all configurations ($m=0,1,2,3$) exhibit the expected decay in electric energy, their behavior with respect to physical invariants differs significantly. Including the constant function $v\mapsto 1$ for $m=1,2,3$ reduces the relative error in mass from approximately $10^{-9}$ to $5\cdot 10^{-11}$. The absolute error of the momentum decreases from approximately $10^{-7}$ to $10^{-11}$ when the function $v\mapsto v$ is included ($m=2,3$). Consequently, both quantities can be considered numerically conserved for $m\geq 2$. As argued in [10] the total energy is not expected to be conserved; however, including $v\mapsto v^{2}$ reduces the error by approximately two orders of magnitude for $m=3$.
@@ 4.3 Mesh- and rank-adaptive transport problem
The initial discretization is based on a regular triangulation with $n_{x}=2\cdot 16^{2}$ and $n_{v}=2\cdot 32^{2}$ elements and a time step of $\tau=5\cdot 10^{-3}$. We estimate the error by projecting onto discontinuous linear finite elements, as described in section 3.4. An element is refined if the error exceeds $10^{-3}$ and all children of a parent element are coarsened if the sum of their error indicators falls below $0.15\cdot 10^{-3}$. The rank adaptive integrator is applied with a truncation tolerance of $10^{-4}$.
@@ 5 Conclusion
We developed and analyzed a conservative DG discretization of the DLRA approach for the numerical solution of the Vlasov–Poisson equation. We presented the conservative integrator from [7, 10], which is based on fixing some velocity basis functions, in a modified form that avoids normal equations in the L-step and suggests a projector splitting scheme. We rigorously derived the DG formulation for the DLRA equations in terms of Friedrichs’ systems, including appropriate numerical fluxes and discrete differentiation operators. Combined with the unconventional rank-adaptive integrator, we proved the conservation of mass and momentum in our numerical solver. The numerical experiments demonstrate the feasibility of our approach, confirm the conservation of physical invariants, and highlight the possibility for mesh adaptivity.
Despite these advances, several questions remain open and could be addressed in future work. First, extending the method to non-periodic boundary conditions requires additional considerations. Second, more sophisticated error estimators and mesh adaptive strategies need to be designed for the DG solvers to maintain the conservation of invariant quantities. Also, higher-order time-stepping schemes, such as those suggested in [20], could be considered. So far, we have established conservation properties only for the unconventional integrator, but certain quantities may also be conserved under the projector splitting approach, which might merit a more detailed investigation.

### v2

@@ 3.3 Conservation of mass and momentum
$\Big{(}\big{(}\hat{F}(f^{n}),1\big{)}_{v},\psi\Big{)}_{x}$ $=\sum_{s}-\Big{(}\hat{\mathrm{d}}_{x_{s}}\big{(}v_{s},f\big{)}_{v},\psi\Big{)}_{x}+\Big{(}E_{s}\big{(}\hat{\mathrm{d}}_{v_{s}}f^{n},1)_{v},\psi\Big{)}_{x}$
for all $\psi\in\mathcal{V}_{h,p}^{(x)}$. Lemma 3.1 implies
$\big{(}\hat{\mathrm{d}}_{v_{s}}f^{n},1\big{)}_{v}=0.$
Applying the same lemma to the first term gives
$\Big{(}\big{(}\hat{F}(f^{n}),1\big{)}_{v},\psi\Big{)}_{x}$ $=\sum_{s}\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}j_{s}\partial_{x_{s}}\psi\,\mathrm{d}\bm{x}-\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}n_{s}^{(e)}\{j_{s}\}_{e}\,[\psi]_{e}\,\mathrm{d}S,$
thereby showing (44). Letting $\psi=1\in\mathcal{V}_{h,p}^{(x)}$ in (44) demonstrates conservation of total mass. ∎
@@
$\frac{1}{\tau}(j_{s}^{n+1}-j_{s}^{n},\psi)_{x}+\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}-\bm{\sigma}_{s}^{n}\cdot\nabla\psi+E^{n}_{s}\rho^{n}\psi\,\mathrm{d}\bm{x}+\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}\big{(}\bm{n}^{(e)}\cdot\{\bm{\sigma}_{s}^{n}\}_{e}\big{)}\,[\psi]_{e}\,\mathrm{d}S=0$ (45)
for all $\psi\in\mathcal{V}_{h,p}^{(x)}$. This corresponds to a DG discretization of the continuity equations (38) using central fluxes and explicit Euler time stepping as in (32). Hence, the discrete local continuity equation holds for the momentum density and the total momentum $\int_{\Omega^{(x)}}\bm{j}^{n}\,\mathrm{d}\bm{x}$ is conserved.
@@
Letting $U=v_{s}$, so that $j_{s}=\varphi_{U}$, and applying Lemma 3.3, we obtain
$j_{s}^{n+1}=j_{s}^{n}+\tau\Big{(}\big{(}\hat{F}(f^{n}),v_{s}\big{)}_{v},\psi\Big{)}_{x}\,$,
with
$\Big{(}\big{(}\hat{F}(f^{n}),v_{s}\big{)}_{v},\psi\Big{)}_{x}$ $=\sum_{s^{\prime}}-\Big{(}v_{s^{\prime}}\,\hat{\mathrm{d}}_{x_{s^{\prime}}}\big{(}v_{s},f\big{)}_{v},\psi\Big{)}_{x}+\Big{(}E^{n}_{s^{\prime}}\big{(}\hat{\mathrm{d}}_{v_{s^{\prime}}}f^{n},v_{s})_{v},\psi\Big{)}_{x}$
for all $\psi\in\mathcal{V}_{h,p}^{(x)}$. In this case Lemma 3.1 yields
$\big{(}\hat{\mathrm{d}}_{v_{s}^{\prime}}f^{n},v_{s}\big{)}_{v}=\delta_{s^{\prime},s}\rho^{n}.$
Applying the same lemma also to the first term hence gives
$\Big{(}\big{(}\hat{F}(f^{n}),v_{s}\big{)}_{v},\psi\Big{)}_{x}=\sum_{s}$ $\sum_{T\in\mathcal{T}_{h}^{(x)}}\int_{T}\sum_{s^{\prime}}\sigma_{s,s^{\prime}}\partial_{x_{s}}\psi\,\mathrm{d}\bm{x}-\sum_{e\in\mathcal{F}_{h}^{(x)}}\int_{e}n_{s}^{(e)}\big{\{}\sum_{s^{\prime}}\sigma_{s,s^{\prime}}\big{\}}_{e}\,[\psi]_{e}\,\mathrm{d}S$
$+\big{(}E^{n}_{s}\rho^{n},\psi\big{)}_{x}\,$,
thereby showing (45). For the conservation of momentum, we note that in the continuous setting one has [12]
@@
This also holds for the discrete version, since the computed $\bm{E}^{n}$ is the exact electric field corresponding to the (discrete) density $\rho^{n}$; see Section 3.2. Setting $\psi=1\in\mathcal{V}_{h,p}^{(x)}$ in (45), and using the fact that $\int_{\Omega^{(x)}}\bm{E}\,\mathrm{d}\bm{x}=0$, we conclude that the total momentum is conserved. ∎
We remark that the key observation for deriving the local continuity equations–and thus the associated conservation properties–is that the updates for $\bm{K}$ and $S$ can be expressed using $\hat{F}$ (Lemma 3.2). However, this property holds only for the central flux. For the upwind flux additional jump terms appear and the local continuity equations in the presented form are hence not satisfied when using the upwind flux. Depending on the basis $\tilde{\bm{X}}^{n+1}$, physical quantities may still be conserved globally. It remains to be investigated whether the algorithm can be modified to also conserve these quantities when employing the upwind flux.
@@ 3.4 Spatial adaptivity
We consider the general setup of section 3 (which applies to both the K- and L-step) and employ an elementwise error indicator based on the projection of a DG solution. Specifically, given a DG function $\bm{U}\in(V_{h,p}^{(y)})^{r}$ of polynomial order $p$, we project it onto a DG space of order p1p-1 within each element $T\in\mathcal{T}_{h}^{(y)}$, resulting in a function $\tilde{\bm{U}}\in V_{h,p-1}^{(y)}$ that satisfies
[…]
@@
In order to remove unnecessary refinements and maintain computational efficiency, a coarsening procedure is applied. The same error indicator is used to assess whether refinement is still required. Specifically, all children of a parent element are removed if the sum of their error indicators falls below $c\,\epsilon$, where $0<c<1$ is a safety factor.
@@ 4.1 Landau damping in one spatial dimension
The results for the conservative unconventional integrator with truncation to a fixed rank of $10$, shown in Figure 1, closely resemble those in [12, Fig. 1]. While all configurations ($m=0,1,2,3$) exhibit the expected decay in electric energy, their behavior with respect to physical invariants differs significantly. Including the constant function $v\mapsto 1$ for $m=1,2,3$ reduces the relative error in mass from approximately $10^{-9}$ to $5\cdot 10^{-11}$. The absolute error of the momentum decreases from approximately $10^{-7}$ to $10^{-11}$ when the function $v\mapsto v$ is included ($m=2,3$). Consequently, both quantities can be considered numerically conserved for $m\geq 2$. As argued in [12] the total energy is not expected to be conserved; however, including $v\mapsto v^{2}$ reduces the error by approximately two orders of magnitude for $m=3$.
@@ 4.3 Mesh- and rank-adaptive transport problem
The initial discretization is based on a regular triangulation with $n_{x}=2\cdot 16^{2}$ and $n_{v}=2\cdot 32^{2}$ elements and a time step of $\tau=5\cdot 10^{-3}$. We estimate the error by projecting onto discontinuous linear finite elements, as described in section 3.4. An element is refined if the error exceeds $10^{-3}$ and all children of a parent element are coarsened if the sum of their error indicators falls below $0.15\cdot 10^{-3}$. The rank adaptive integrator is applied with a truncation tolerance of $10^{-4}$. The time step $\tau$ itself was fixed to a small enough value to fulfill CFL conditions at all times and for all meshes. In principle the time step could be adapted to the grid or substepping could be applied, but this was not implemented.
@@ 5 Conclusion
We developed and analyzed a conservative DG discretization of the DLRA approach for the numerical solution of the Vlasov–Poisson equation. We presented the conservative integrator from [8, 12, 10], which is based on fixing some velocity basis functions, in a modified form that avoids normal equations in the L-step and suggests a projector splitting scheme. We rigorously derived the DG formulation for the DLRA equations in terms of Friedrichs’ systems, including appropriate numerical fluxes and discrete differentiation operators. Combined with the unconventional rank-adaptive integrator, we proved the conservation of mass and momentum in our numerical solver. The numerical experiments demonstrate the feasibility of our approach, confirm the conservation of physical invariants, and highlight the possibility for mesh adaptivity.
Despite these advances, several questions remain open and could be addressed in future work. First, extending the method to non-periodic domains and corresponding boundary conditions requires additional considerations, potentially based on the ideas in [23]. Second, more sophisticated error estimators and mesh adaptive strategies need to be designed for the DG solvers to maintain the conservation of invariant quantities. Also, higher-order time-stepping schemes, such as those suggested in [22], could be considered. So far, we have established conservation properties only for the unconventional integrator, but certain quantities may also be conserved under the projector splitting approach, which might merit a more detailed investigation.